17. Z Co. has total assets of $150,000, total stockholders equity of $80,000, total annual revenues of $460,000, cost of gooCould you solve clearly? Thanks a lot!!
Answer #1

Total assets = Total Liabilities + Total stockholders equity

150,000 = Total Liabilities + 80,000

Total Liabilities = 150,000 - 80,000 = 70,000

Percentage of total liabilities

= 70,000/150,000

= 47%